### v1

# Beyond Baby Blues: The Child Penalty in Mental Health in Switzerland

Abstract
This paper investigates the mental health penalty for women after childbirth in Switzerland. Leveraging insurance data, we employ a staggered difference-in-difference research design. The findings reveal a substantial mental health penalty for women following the birth of their first child. Approximately four years after childbirth, there is a one percentage point (p.p.) increase in antidepressant prescriptions, representing a 50% increase compared to pre-birth levels. This increase rises to 1.7 p.p. (a 70% increase) six years postpartum. The mental health penalty is likely not only a direct consequence of giving birth but also a consequence of the changed life circumstances and time constraints that accompany it, as the penalty is rising over time and is higher for women who are employed before childbirth.
JEL classification: J13, J16, I10
Keywords: Parenthood, fertility, depression, antidepressants, parental leave, staggered difference-in-difference
1 Introduction
Mental health awareness has increased in recent years and is key to welfare in a society (Votruba:2016). Approximately 13% of women and 8% of men in Switzerland between 15 and 44 years experienced a mental health crisis that needed care in 2023, and the costs for mental illness in Switzerland are estimated to be around 7 billion Swiss francs per year (SRF:2023a). [^1] This is around 2% of the Swiss yearly government spending. In general, the risk of suffering from mental health problems varies between individuals due to environmental, social and economic risk factors and individual attributes and behaviours (WHO:2012). [^2] Environmental factors include, for example, cultural beliefs, economic policies or gender inequality. Social and economic factors include, for example, the engagement with family members or friends or earnings. Individual attributes and behaviours include dealing with feelings, managing oneself in life, and participating in social activities. Furthermore, they can be influenced by biological factors, such as genes.
The birth of a child is a significant life event that increases some risk factors for mental health problems. Women carry the physical burden during pregnancy and childbirth, and biological factors change. Afterwards, a child leads to more stress related to sleep deprivation, a change in the family structure and the roles and responsibilities of the parents, career uncertainties and maybe also financial pressure (Heshmati:2023). These circumstances do not only change for women but also for men. Hence, the risk of mental health problems increases. The prevalence of postpartum depression in industrialised countries for women lies between 10% and 15% (WHO:2020) and for men around 5% (Bradley:2011). This higher risk of psychiatric disorders does also not decrease if the pregnancy was planned (Munk:2015). Hence, even if a couple wants a child, it does not prevent parents from suffering from a mental health crisis.
Several studies summarised by Blanchflower:2009, Clark:2008 and Dolan:2008 look at the association between parenthood and self-reported well-being and find a negative effect between having a child and happiness in the short-run. However, self-reported outcome variables might be prone to systematic under-reporting or selection bias due to lower response rates from individuals with health problems (Baetschmann:2016). Recent research by Ahammer:2023 has shed light on the different effects of having a child on mental health among men and women using drug prescription data. Women in Austria and Denmark face enduring mental health challenges after giving birth. The consumption of antidepressants significantly increases post-childbirth. Similarly, the percentage of men taking antidepressants also increases after childbirth. However, the increase is higher for women, with no significant reduction in the mental health gap between men and women over the years. This gap appears to be less related to the act of childbirth but more to the evolving roles and responsibilities that follow (Ahammer:2023).
Women generally invest more time in childcare and household duties than men (Borra:2021; Guryan:2008). Consequently, they are more likely to reduce their working hours outside the home (e.g. Kleven:2020). Even when women outearn their husbands, they spend more time doing household chores (Bertrand:2015). Hence, the evolving roles and responsibilities that mothers face both at home and in the workplace can significantly impact their mental health as time pressure mediates the relationship between parenthood and mental health (Ruppanner:2019). The combination of these dual roles, along with insufficient support at home, often leads to stress and sleep deprivation, thereby increasing the risk of mental health issues for women more than for men, resulting in an increased mental health gender gap.
Simple measures such as household and childcare assistance, extended time off work, or access to childcare facilities can significantly reduce the increasing mental health burden after childbirth. Therefore, family policies play a crucial role in reducing the risk of mental health issues such as depression, posttraumatic stress disorder, anxiety, and postpartum psychosis (e.g., Chatterji:2008; Avendano:2015; Aitken:2015; Hewitt:2017). A systematic review study by Heshmati:2023 suggest that an introduction of maternity leave improves mental health outcomes after childbirth as found in the majority of studies (e.g. Chatterji:2008; Avendano:2015) but the length of the leave seems to be crucial. Studies in Austria find that an increased maternity leave exacerbates the gender gap in mental health through further enforcement of traditional gender roles (e.g., Chuard:2023; Ahammer:2023). [^3] The parental leave reform extended parental leave from 24 months up to 36 months. The maximum leave for one parent was increased from 18 months to 30 months. The evidence on the effect of paternity leave on maternal mental health is scarce. Some research suggests that women’s mental health improves when partners have access to paternity leave (Bilgrami:2020) or to flexible paternity leave policies (Persson:2019).
This study primarily contributes to the literature on the effect of a child on the mental health of women using anonymous prescription data and looking at medium-term outcomes (Ahammer:2023; Kravdal:2017). [^4] Ahammer:2023 use data from Austria and Denmark and Kravdal:2017 use data from Norway. We investigate the mental health penalty for women after the birth of their first child in Switzerland by using insurance data from two of the biggest Swiss insurance companies and employing a staggered difference-in-difference estimation strategy (Callaway:2021). Mental health is proxied by antidepressant prescriptions as the primary outcome. Additionally, the number of visits to psychiatrists and the number of general practitioners (GP) are used as secondary measures, as these healthcare providers issue most of these prescriptions (NZZ:2019). Switzerland has one of the highest female employment rates (BfS:2023c) in Europe, but many mothers only work part-time. This situation may be influenced by the country’s relatively short maternity leave (16 weeks), the absence of paternity leave until 2021 (extended to two weeks in 2021), and the high cost of childcare. Furthermore, Switzerland has one of the highest gender pay gaps in Europe (BfS:2023c), which may also contribute to the tendency for women to reduce their working hours more than men. As a result, many women in Switzerland face a double workload without the same level of support found in more family-friendly countries like Denmark or Norway, which may lead to different outcomes for working mothers.
The findings reveal that four years after childbirth, the percentage of women having an antidepressant prescription in a specific month rises from 2% to 3%, representing a 50% increase. [^5] Women who are about to have a child, especially in the years leading up to childbirth, are less likely to have an antidepressant prescription compared to women who will never have a child or those who already have children. This percentage further climbs to 3.3% six years post-childbirth, marking a 75% increase. The results align with the findings of Ahammer:2023 for Austria and Denmark. The effect found is substantial and economically relevant as mental illnesses impose high costs on the healthcare system (Swissinfo:2008). Additionally, considering that mental health is a prerequisite for labour market participation, there is a notable loss of productivity in an economy already experiencing a shortage of skilled workers (e.g. Biasi:2021). While we do not find significant effects for visits to psychiatrists or GPs in our main specification, we observe an increase in psychiatrist visits in certain alternative specifications, suggesting some sensitivity of the results to sample choice.
Additionally, this paper sheds light on some potential mechanisms by looking at heterogeneous treatment effects. There is a higher penalty for women employed before childbirth than those not employed. [^6] In Switzerland, accident insurance is paid by the employer if an individual is employed. Accident coverage included in the health insurance is only needed for individuals who are unemployed, out of the labour market or self-employed. Therefore, we can identify the employment status. The double workload at home and work outside the home might exacerbate the mental health penalty. Moreover, despite some evidence from the literature which finds that cesarean sections are associated with a higher risk of postpartum depression (Moameri:2019; Xu:2017), there is no indication of a difference in antidepressant prescriptions by the mode of delivery. Hence, the surgery itself does not seem to be the reason for the mental health penalty. Similarly, despite some literature showing that individuals with low socioeconomic status are more likely to suffer from mental health problems after childbirth (e.g. Goyal:2010; Rich:2006), we find that the mental health penalty is lower for individuals with a low income. This can partly be explained by the fact that those women are more likely to not be employed, as income correlates with employment. We also find that the number of first-time antidepressant prescriptions is stable in the first six years after childbirth, which suggests that the increase in antidepressant prescriptions does not only stem from women having postpartum depression and then getting addicted to the drugs. Concluding, these different findings suggest that the increase in antidepressant prescriptions only partly comes from postpartum depression. The other part of the increase probably stems from changing circumstances, leading to more stress.
Last, this paper contributes to the literature on the effect of paternity leave on maternal mental health. We investigate the effect of the introduction of paternity leave in 2021 in Switzerland on the mental health of the mother. The literature on this highly relevant question is scarce. Mental health is crucial for women to enter the labour market after childbirth (e.g., Biasi:2021), and as mentioned, mental illness is costly for an economy. Therefore, economists should also consider the impact of family policies on mental health outcomes and not only on labour market outcomes (e.g., Kleven:2019; Olivetti:2017; Havnes:2011). Since the introduction was in 2021, only short-run effects can be analysed. We do not find any significant effect on mental health. Due to the low power of the analysis - only a few women have an antidepressant prescription in the first two years after childbirth; therefore, the outcome is rare - the null result should be interpreted with caution.
The remainder of the paper is structured as follows: Section 2 explains the healthcare system and provides some facts about women’s general mental health, the labour force participation and the different family policies in Switzerland. Section 3 describes the data used, how the sample has been selected for the analysis and shows some descriptives. Section LABEL:stagg_diff_diff:_empirical_strategy explains the empirical strategy, and Section LABEL:results_main presents the different results, including robustness checks, results for men and heterogeneities. The results are further discussed in Section LABEL:discussion, and potential mechanisms are explained. Last, Section LABEL:1_conclusion concludes. Additional descriptives and results are presented in Appendix LABEL:appendix:_staggered_diff_diff. Appendix LABEL:appendix:_effect_on_men presents the effect of childbirth on men’s mental health, while Appendix LABEL:appendix:_paternal_leave covers the impact of paternal leave on mothers’ mental health.
2 Institutional Setting
2.1 Healthcare System
Having basic health insurance is mandatory in Switzerland. Residents of Switzerland can choose among 50 different health insurers, but over 90% of individuals are insured at the ten biggest health insurances (SRF:2024). The contracts are made for one year, meaning individuals can change insurers yearly. Between 2010 and 2022, around 10% of individuals regularly changed health insurers (versicherung_schweiz:2021). In the event of illness, maternity and accident, the health insurance protects the insured people, and the insurer is not allowed to treat insured individuals differently, for example, by differentiating between healthy and non-healthy individuals. This also applies to insurance admission. The insurance is financed by the monthly contributions from the individuals (premium), with the individuals’ cost-sharing (deductible, excess and hospital contributions) and partly with contributions from the cantons and the Swiss Confederation. How much an individual pays monthly depends on the place of residence and the chosen form of insurance, for example, how high the deductible is and what the first point of contact is in case of medical need (e.g. a GP, a telemedicine service or directly a specialist). The average monthly premium for an adult is slightly over CHF 300. The most often used yearly deductible is CHF 2,500 (BAG:2023). Until the deductible is met, an individual pays for the medical treatments by himself or herself. Afterwards, the insurance covers the costs. Despite paying the costs individually, all services are billed to the insurance company, which reclaims the costs from the individuals. Hence, the insurance company knows how much of the deductible has already been used and what services and treatments an individual has used in a year.
Additionally, individuals in Switzerland can sign up for supplementary insurance, which covers additional services, such as alternative medicine, glasses, or a private room in the hospital. It is not mandatory to have supplementary insurance, the insurance companies can refuse individuals, and it is possible to have basic and supplementary insurance at different insurance companies. Therefore, supplementary insurance is not included in the main analysis for consistency reasons. [^7] Data on the private hospital insurance is used in the identification section to explore if individuals are signing up for it at a certain point before having a child, which could be a sign of planning a child.
2.2 Mental Health
Around 9% of women suffer from a major depression in Switzerland (BfS:2023a) and around 5% to 7% of the Swiss population take antidepressants at least once a year (BfS:2019). The prevalence is lower for younger women, as the risk of depression increases with age. Furthermore, it is noteworthy that many antidepressants are prescribed by GPs rather than psychiatrists, highlighting the crucial role GPs play in the mental health care of the Swiss population (Obsan:2022). The basic insurance covers the costs of antidepressants and psychiatrist visits (after using up the deductible). From 2022 on, it also covers visits to a psychotherapist who is not working for a psychiatrist (BAG:2022). For consistency reasons, those visits are not included in the analysis.
2.3 Family Policies and Labour Force
As pointed out in the literature, part of the mental health problems after childbirth are likely mediated by an increased level of stress, sleep deprivation and workload due to household responsibilities, childcare duties, and work outside the home. Therefore, family policies are crucial in supporting maternal well-being, primarily if the mother works outside the home.
The employment rate for women in Switzerland lies around 90% before childbirth (BfS:2022). This is a high participation rate compared to other European countries. The percentage of women working after a first child decreases to 80% and after a second to 69%. As their children grow older, more women return to work, with about 87.5% of mothers with children aged 12 to 14 working again. Most mothers work only part-time, namely 78.1 %, while only 18% of Swiss fathers do (BfS:2022). Part of the reason for part-time work is that childcare is expensive in Switzerland (BSV:2024a). In 2003, the Swiss government decided to incentivise the building of new childcare places, which was relatively successful in increasing the availability but not in reducing childcare costs.
Maternity and paternity leave can help to reduce the mental health burden of mothers after childbirth. In 2005, a 16-weeks maternity leave has been introduced in Switzerland (BSV:2024). Compared to other European countries, this is a relatively short period. A two-week paternity leave has only been introduced in 2021 (BSV:2024). Furthermore, Switzerland has no flexible parental leave system that allows parents to choose who stays home with the children.
3 Data
3.1 Data Sources
This paper uses data from two of the biggest Swiss insurance companies (CSS and Swica), including all incurred insurance costs for each individual over time. It is a selected sample of women insured at CSS and Swica, which corresponds to roughly 25% of the Swiss population. The dataset from CSS contains information on insured women and men born between 1971 and 2003 from the years 2010 to 2022, with information on the drugs that a person buys and the visits to GPs and psychiatrists over the years. The data from Swica contains the same information but only for women and starts in 2012. Hence, all costs related to mental health that occur for an individual are observed. Additionally, the dataset provides each insured person’s basic sociodemographic characteristics, such as age, civil status, and region of residence.
The advantage of using insurance data is that there are no self-reported health outcomes, which could easily lead to systematic under-reporting (Bound:2001). Furthermore, survey data could suffer from selection bias due to lower response rates from individuals with health problems. [^8] Theoretically, some individuals might prefer to pay for antidepressants out of pocket to prevent their insurance from knowing about their mental health issues, potentially leading to an under-reporting bias. However, this scenario is highly improbable for several reasons. First, personal health data is protected by law, ensuring that insurance providers cannot disclose an individual’s use of antidepressants. Second, individuals can switch basic insurance plans annually, with other insurers prohibited from refusing coverage. Third, when applying for supplementary insurance, a mandatory health assessment requires disclosing all relevant health information. Fourth, if a psychiatrist or GP directly distributes the antidepressants, the cost is included in the overall bill for the visit. Even if dispensed by a pharmacy, individuals must explicitly request to pay out of pocket. A disadvantage is that mental health changes are only observed if the person uses some medical treatment, i.e. either some drugs or a visit to a doctor. Hence, small changes in the mental health conditions of individuals are not detectable. Furthermore, due to the sensitivity of the data, it cannot be linked to other administrative sources. Since individuals can easily switch insurance companies yearly, the dataset only provides a partial panel for some women.
Additionally, this paper uses administrative data from the Swiss government, i.e. income compensation data, population and household statistics, vital statistics and social security data. The data is used to check if the insurance data is representative of the Swiss population. For more information on the representativeness of the data, see Appendix LABEL:appendix:_data_descriptives.
3.2 Selected Sample
The analysis focuses on women born between 1985 and 1995. We exclude older women from the sample because they are more likely to have had their first child before 2010. For example, women born in 1985 would have been 25 years old in 2010, and while many may have delayed childbearing, some are likely to have become mothers before 2010. Hence, second births would be misclassified as first births. [^9] Some births may still be misclassified, with a slightly higher potential for misclassification among individuals insured with Swica, as the data only begins in 2012 rather than 2010. Conversely, women born after 1995 are less likely to have had their first child by 2022, as those born in 1995 would only be 27 years old. Consequently, including women born after 1995 would shift the distribution of the age at the first birth to the left. Therefore, they are excluded from the analysis. [^10] These thresholds have been defined with the median age of women giving birth to their first child in Switzerland in mind, which is 31 years (BfS:2024). The sample is further refined to include only women who gave birth to their first child between 20 and 40.111111By construction, the oldest women in the sample are 37 years old when they have their first child - women born 1985 that have their first child in 2022. Hence, only the lower age is binding. This age range allows us to focus on women’s reproductive years.
We further constrain the samples to the years between 2012 and 2022, a choice influenced by the introduction of the DRG on the 1st of January 2012 (DRG:2024). [^12] A Diagnostic-Related Group (DRG) is a system used by health insurance companies to categorise hospitalisation costs and determine payment amounts for hospital stays. Instead of billing for each individual service received, a fixed payment amount is predetermined based on the DRG assigned to your stay.. This significant event in the healthcare system could affect data reliability prior to 2012. After applying these criteria, the sample of women comprises $N=83,841$ individuals.
To further mitigate issues regarding misclassifying first births, we use a balanced sample of women who have remained with the same insurance provider between 2012 and 2022. While this approach helps reduce the risk of bias caused by compositional changes over time, as it captures the same group of women throughout the study period, it may introduce a limitation by focusing on a specific population, as the individuals have not changed insurance for over a decade. [^13] Because women have a child at different points in time, and we only contemplate the dataset from 2012 to 2022, we still have a compositional change if using a staggered difference-in-difference approach (Callaway:2021) but a smaller one than by using an unbalanced sample. Assuming a woman has her first child in 2018, we contemplate six years before and five years after childbirth. Hence, if we look at the effect six years after childbirth, this woman is not in the sample anymore. To eliminate all compositional changes, we would have to restrict the sample to women for whom we contemplate the whole event-time of interest. This would reduce the sample size even more. Table 3.3.2 in Section 3.3.2 shows that the balanced sample only differs from the unbalanced sample in terms of nationality, language, and number of children. Thus, the balanced sample appears mostly representative of the unbalanced sample and, hence, of the broader Swiss population.
This decision to use a balanced sample comes at the cost of reducing the sample size of women to $N=25,120$. Hence, the power of the analysis is reduced. An unbalanced sample is used to check for robustness. Furthermore, a balanced sample only going until 2019 ($N=18,714$) is used to avoid the potential bias of the COVID-19 pandemic on mothers’ mental health. If mothers with specific mental health choose to have a child during the pandemic and, at the same time, the pandemic has an impact on antidepressant prescriptions, this could bias the results in both directions. [^14] We do not find any change in the pattern of antidepressant prescriptions due to the Corona pandemic (see Figure LABEL:fig:descriptives_antidepri_over_years in Appendix LABEL:1_data_paternity_leave).
3.3 Descriptive Statistics
3.3.1 Outcome Variables
The primary outcome variable is the mental health status of a woman, proxied by whether a woman has an antidepressant prescription (ATC Code: N06A) in a particular month. Since antidepressant prescriptions can last several months, the outcome variable is smoothed over the months the individual likely consumed the antidepressant. This smoothing is based on the number of prescribed packages and pills per package. Typically, patients consume one pill per day, so a package containing 30 pills is assumed to last one month. Additionally, if a woman lacks an antidepressant prescription for one or two months between two months with prescriptions, continuous consumption during these months is assumed. This smoothing approach provides a less volatile outcome variable without altering the results. [^15] A robustness check with the non-smoothed outcome can be found in Appendix LABEL:appendix_non_smoothed. Nevertheless, antidepressant prescriptions are potentially still a noisy proxy for mental health, and they most likely underestimate mental health problems, as they are only prescribed in severe cases.
In addition to the primary outcome, we consider other mental health proxies, such as the number of visits to a GP or psychiatrist. These measures are relevant since GPs or psychiatrists prescribe most antidepressants. It is essential to note that there are limitations to these outcomes. Increased GP visits may result from various factors unrelated to mental health, such as childcare-related illnesses, and time constraints, particularly in the first years after childbirth, might influence the number of visits to a psychiatrist and GP more than the antidepressant prescriptions. Furthermore, many GPs are located in a Health Maintenance Organisation (HMO) with other physicians, such as gynaecologists or specialists. We combine GP and HMO visits. Therefore, there is a slight measurement error in this outcome variable. Furthermore, the number of visits to the GP or HMO can only be identified consistently from 2012 to 2022 for the CSS data, which leads to a smaller sample size for this outcome. The outcome variables are measured relative to the event-time of birth $t$, with $t+1,t+2,\dots$ representing the months post-childbirth, and $t-1,t-2,\dots$ the months pre-childbirth.
Figure 1 illustrates the number of women with an antidepressant prescription in the months before and after childbirth. The red-shaded area represents the pregnancy period, culminating in the first child’s birth at $t=0$. Approximately 2% of women have an antidepressant prescription in any given month before childbirth. [^16] If we look at yearly prescriptions, approximately 4% of women have an antidepressant prescription before childbirth. This number is slightly lower than the average in the Swiss population as antidepressant consumption increases with age. This number increases over the years following childbirth, reaching 2.5% four and 3% six years after childbirth. The observed decrease in antidepressant prescriptions during pregnancy and the first year postpartum is likely due to the recommendation that it is better for the child’s health if the mother avoids consuming antidepressants.
Figure 1: Descriptives: Antidepressants prescriptions
Note: The figure shows the percentage of women taking antidepressants around the time of birth ($t=0$: time of first birth). The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=25,120$)
To assess whether the increase in antidepressant use for women is due to an immediate rise after childbirth leading to drug dependence, we examine first-time antidepressant prescriptions. Since the prescription data only begins in 2012, identifying actual first-time prescriptions is challenging. As we move further in time, the likelihood increases that the observed prescriptions are genuinely first-time prescriptions, explaining the negative trend in antidepressant use among women without children (blue line) in Figure 2. To minimise the problem, we select women with at least one year without antidepressant prescriptions before their first recorded prescription. The number of women receiving their first antidepressant prescription stabilises at a consistent level in the years after childbirth. This suggests that the observed increase in antidepressant use is not only explainable by postpartum depression.
Figure 2: Descriptives: First-time antidepressants prescriptions
Note: The figure shows the percentage of women taking for the first time antidepressants around the time of first birth ($t=0$). In the sample are only women who have at least one antidepressant prescription between 2012 and 2022 and who have had at least one year without antidepressant prescriptions before their first recorded prescription. The pink line shows the linear trend before and after childbirth for women giving birth in $t=0$. The blue line shows the linear trend for women who never have a child but get assigned random birthdates. The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=7,489$)
Figure 3 shows that a woman visits on average 0.12 times per month a psychiatrist. [^17] Slightly more than 3.5% of women go to a psychiatrist before childbirth. The number of visits decreases during pregnancy, childbirth and the first years after birth, likely due to time constraints. Two years after childbirth, the number of visits increases, reaching approximately 0.14 times per month six years post-childbirth.
Figure 3: Descriptives: Visits to psychiatrist
Note: The figure shows the number of visits to a psychiatrist around the time of birth ($t=0$: time of first birth). The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=25,120$)
Figure 4 depicts that, on average, women visit a GP or HMO around 0.18 times per month (over two visits per year). The frequency of visits increases during pregnancy, likely due to gynaecological check-ups at HMO centres, then declines after childbirth. However, the rate of visits begins to rise again, reaching 0.22 visits per month by the time children are six years old.
Figure 4: Descriptives: Visits to GP or HMO
Note: The figure shows the number of visits to a GP or HMO around the time of birth ($t=0$: time of first birth). The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=16,368$)
3.3.2 Covariates
We control for the individual’s age at their first birth in the analysis and use additional covariates to analyse heterogeneities. Controlling for age is important, as the risk of depression increases with age and because women who are mentally unhealthy might postpone having a child. Hence, women having a child at different ages may have different mental health trends. The average age of having a first child is around 29 years, and the age distribution is shown in Figure 5.181818The average age of Swiss women when they give birth for the first time is 30. However, the mean in our sample is lower due to selecting only women born between 1985 and 1995 and having the years from 2010 to 2022. By construction, the highest age at the birth of a first child is 37 years, and the lowest is 20 due to some sample restrictions.
Figure 5: Descriptives: Age at first birth
Note: The figure shows the age distribution of women having a child between 2012 and 2022. ($N=25,120$)
Accident coverage is a proxy for employment status, as it is typically needed by those unemployed, out of the labour force, or self-employed. At the same time, employers generally provide it for their employees. Figure 6 shows a notable increase in accident coverage following childbirth, suggesting a rise in non-employment among women. Around 80% of women do not have accident coverage before childbirth, implying an 80% employment rate. [^19] The decrease in accident coverage pre-birth is partly due to young women and men not yet having accident insurance because they are, for example, studying or doing an apprenticeship. This contrasts with the 90% pre-childbirth employment rate for women in Switzerland as reported by official statistics, indicating a discrepancy due to self-employment (BfS:2022). 8.1% of women without and 11.8% of women with children in Switzerland are self-employed. As depicted by Figure 6, accident coverage rises to about 30% (70% employed) post-childbirth, and further increases to 40%, suggesting 60% employment. Official figures show a decrease in employment to 80% after the first child and 69% after the second (BfS:2022). Thus, the proxy continuously underestimates employment by approximately ten percentage points (p.p.). This ten p.p. discrepancy is likely due to the exclusion of self-employed individuals. [^20] There is a concern that individuals might forget to cancel their accident insurance when they find a job, as it is inexpensive. However, the more common scenario is that people stop working after childbirth and need to sign up for accident coverage. Since accident insurance is mandatory and reminders are sent if it is not in place, this issue is less pronounced.
Figure 6: Descriptives: Accident coverage
Note: The figure shows the proportion of women having accident coverage around the time of birth ($t=0$: time of first birth). The red shaded area illustrates the pregnancy, and the shaded area represents the 95% confidence interval. ($N=25,120$)
The left part of Table 3.3.2 shows the sociodemographic characteristics of women in the month they have a child, separated by the different samples used, whereas the right part of the table shows it separated by women having an antidepressant prescription or not.
In the balanced sample, 19% of women receive subsidies for health care, 63% of women have more than one child, and 30% have a cesarean section. Furthermore, 78% of the women have German as a contact language, 19% have French, and 3% have Italian. The balanced and the unbalanced samples differ in some sociodemographic characteristics. For example, the balanced sample has a higher number of Swiss women and fewer women who speak English. A reason could be that some foreign women entered Switzerland later than 2012 and, hence, are not insured at the same insurance for the whole period. Similarly, some sociodemographic characteristics differ between the sample going until 2022 and the sample only going until 2019. On average, the latter has a lower age at first birth. By construction, the oldest woman giving birth can only be 34 years old instead of 37, which pushes the average age downwards. The unbalanced has more women with only one child by construction. We do not see all the years for the women in the unbalanced sample, hence, we miss some second births.
Comparing women who have an antidepressant prescription with those who do not, we see that slightly more women speaking French and fewer women speaking German have an antidepressant prescription. This pattern can also be seen in the regions of residence, as more women living in the French-speaking part of Switzerland have an antidepressant prescription compared to, for example, central Switzerland. As expected, women with an antidepressant prescription are more likely to have an accident insurance, meaning that they are more likely to not be employed. Furthermore, they are more likely to have a subsidy for health care, have a cesarean section and only have one child.
Table 1: Descriptives: Comparison of different samples for staggered difference-in-difference analysis

### v2

@@
 This increase rises to 1.7 p.p.
-(a 70% increase) six years postpartum.
-The mental health penalty is likely not only a direct consequence of giving birth but also a consequence of the changed life circumstances and time constraints that accompany it, as the penalty is rising over time and is higher for women who are employed before childbirth.
+(a 75% increase) six years postpartum.
+The mental health penalty is likely not only a direct consequence of giving birth but also a consequence of the changed life circumstances and time constraints that accompany it, as the penalty is rising over time and is higher for women who are employed.
 JEL classification: J13, J16, I10
-Keywords: Parenthood, fertility, depression, antidepressants, parental leave, staggered difference-in-difference
+Keywords: antidepressants, depression, fertility, parenthood, staggered difference-in-difference,
 1 Introduction
@@
 Furthermore, they can be influenced by biological factors, such as genes.
-The birth of a child is a significant life event that increases some risk factors for mental health problems.
+The birth of a child is a significant life event that is often associated with happiness.
+However, it also increases some risk factors for mental health problems.
 Women carry the physical burden during pregnancy and childbirth, and biological factors change.
@@
 Hence, even if a couple wants a child, it does not prevent parents from suffering from a mental health crisis.
-Several studies summarised by Blanchflower:2009, Clark:2008 and Dolan:2008 look at the association between parenthood and self-reported well-being and find a negative effect between having a child and happiness in the short-run.
+This paper examines the impact of childbirth on maternal mental health in Switzerland, using anonymized prescription data and a staggered difference-in-differences approach.
+Several studies summarised by Blanchflower:2009, Clark:2008 and Dolan:2008 already looked at the association between parenthood and self-reported well-being and find a negative effect between having a child and happiness in the short-run.
 However, self-reported outcome variables might be prone to systematic under-reporting or selection bias due to lower response rates from individuals with health problems (Baetschmann:2016).
@@
 Similarly, the percentage of men taking antidepressants also increases after childbirth.
-However, the increase is higher for women, with no significant reduction in the mental health gap between men and women over the years.
+However, the increase is higher for women, with no statistically significant reduction in the mental health gap between men and women over the years.
 This gap appears to be less related to the act of childbirth but more to the evolving roles and responsibilities that follow (Ahammer:2023).
+Similar effects of a decline in mental health after childbirth for women are found in Germany (Dehos:2024).
+They further suggest that key factors contributing to this decline may include reduced sleep, less physical activity, and limited leisure time, alongside the extensive involvement in childcare.
 Women generally invest more time in childcare and household duties than men (Borra:2021; Guryan:2008).
@@
 The combination of these dual roles, along with insufficient support at home, often leads to stress and sleep deprivation, thereby increasing the risk of mental health issues for women more than for men, resulting in an increased mental health gender gap.
+[^3] Another strand of the literature shows that intimate partner violance could also be associated with an increase in mental health problems after childbirth as being pregnant and having a child makes women more vulnerable and dependent on their partner (Woolhouse:2012; Chan:2022).
+As we are not able to add any evidence on this, we do not discuss this further.
 Simple measures such as household and childcare assistance, extended time off work, or access to childcare facilities can significantly reduce the increasing mental health burden after childbirth.
@@
 The evidence on the effect of paternity leave on maternal mental health is scarce.
-Some research suggests that women’s mental health improves when partners have access to paternity leave (Bilgrami:2020) or to flexible paternity leave policies (Persson:2019).
-This study primarily contributes to the literature on the effect of a child on the mental health of women using anonymous prescription data and looking at medium-term outcomes (Ahammer:2023; Kravdal:2017).
+Some research suggests that womens’ mental health improves when partners have access to paternity leave (Bilgrami:2020) or to flexible paternity leave policies (Persson:2019).
+This study primarily contributes to the literature on the effect of a child on the mental health of women using anonymized prescription data and looking at medium-term outcomes (Ahammer:2023; Kravdal:2017).
 [^5] Ahammer:2023 use data from Austria and Denmark and Kravdal:2017 use data from Norway.
@@
 Switzerland has one of the highest female employment rates (BfS:2023c) in Europe, but many mothers only work part-time.
-This situation may be influenced by the country’s relatively short maternity leave (16 weeks), the absence of paternity leave until 2021 (extended to two weeks in 2021), and the high cost of childcare.
+This situation may be influenced by the country’s relatively short maternity leave (14 weeks), the absence of paternity leave until 2021 (extended to two weeks in 2021), and the high cost of childcare, as public kindergarten starts when a child is between three and five years depending on the region in Switzerland.
 Furthermore, Switzerland has one of the highest gender pay gaps in Europe (BfS:2023c), which may also contribute to the tendency for women to reduce their working hours more than men.
@@
 Biasi:2021).
-While we do not find significant effects for visits to psychiatrists or GPs in our main specification, we observe an increase in psychiatrist visits in certain alternative specifications, suggesting some sensitivity of the results to sample choice.
+While we do not find statistically significant effects for visits to psychiatrists or GPs in our main specification, we observe an increase in psychiatrist visits in certain alternative specifications, suggesting some sensitivity of the results to sample choice.
 Additionally, this paper sheds light on some potential mechanisms by looking at heterogeneous treatment effects.
@@
 Since the introduction was in 2021, only short-run effects can be analysed.
-We do not find any significant effect on mental health.
-Due to the low power of the analysis - only a few women have an antidepressant prescription in the first two years after childbirth; therefore, the outcome is rare - the null result should be interpreted with caution.
-The remainder of the paper is structured as follows: Section 2 explains the healthcare system and provides some facts about women’s general mental health, the labour force participation and the different family policies in Switzerland.
+We do not find any statistically significant effect on mental health.
+Due to the low power of the analysis, only a few women have an antidepressant prescription in the first two years after childbirth; therefore, the outcome is rare, the null result should be interpreted with caution.
+The remainder of the paper is structured as follows: Section 2 explains the healthcare system and provides some facts about womens’ general mental health, the labour force participation and the different family policies in Switzerland.
 Section 3 describes the data used, how the sample has been selected for the analysis and shows some descriptives.
@@
 Additional descriptives and results are presented in Appendix LABEL:appendix:_staggered_diff_diff.
-Appendix LABEL:appendix:_effect_on_men presents the effect of childbirth on men’s mental health, while Appendix LABEL:appendix:_paternal_leave covers the impact of paternal leave on mothers’ mental health.
+Appendix LABEL:appendix:_effect_on_men presents the effect of childbirth on mens’ mental health, while Appendix LABEL:appendix:_paternal_leave covers the impact of paternal leave on mothers’ mental health.
 2 Institutional Setting
@@
 Hence, the insurance company knows how much of the deductible has already been used and what services and treatments an individual has used in a year.
+Home visits by midwives for postnatal care and monitoring the health of both mother and newborn are covered by basic health insurance within the first 56 days postpartum.
+In cases of premature birth, multiple births, first-time mothers, or cesarean sections, up to 16 visits are permitted.
+Otherwise, the limit is 10 visits.
+During the first 10 days, up to five additional second daily visits are allowed, and any further visits beyond these limits require a physician’s prescription.
+In conclusion, the midwife visits once to twice weekly in the first two months after childbirth (BAG:2024).
 Additionally, individuals in Switzerland can sign up for supplementary insurance, which covers additional services, such as alternative medicine, glasses, or a private room in the hospital.
@@
 For consistency reasons, those visits are not included in the analysis.
+Postpartum depression refers to depressive episodes that arise during pregnancy or within the first year after childbirth.
+Common symptoms include persistent sadness, loss of interest or pleasure in previously enjoyed activities, sleep disturbances, feelings of worthlessness or being a bad mother, frequent crying, and thoughts of death or suicide (APA:2020).
+These symptoms largely mirror those of non-postpartum depression.
+For mild cases of depression, psychotherapy is typically the first line of treatment.
+In moderate cases, psychotherapy is often combined with antidepressant medication, while in severe cases, antidepressant medication is considered essential.
+Depending on individual needs, additional medications such as sleep aids or anxiolytics may also be prescribed (Swissmom:2024).
 2.3 Family Policies and Labour Force
@@
 Maternity and paternity leave can help to reduce the mental health burden of mothers after childbirth.
-In 2005, a 16-weeks maternity leave has been introduced in Switzerland (BSV:2024).
+In 2005, a 14-weeks maternity leave has been introduced in Switzerland.
+During this period, 80% of the salary is paid, and job protection extends for up to 16 weeks (BSV:2024).
 Compared to other European countries, this is a relatively short period.
@@
 The data from Swica contains the same information but only for women and starts in 2012.
-Hence, all costs related to mental health that occur for an individual are observed.
+Hence, all costs related to mental health and childbirth that occur for an individual are observed.
 Additionally, the dataset provides each insured person’s basic sociodemographic characteristics, such as age, civil status, and region of residence.
@@
 [^11] These thresholds have been defined with the median age of women giving birth to their first child in Switzerland in mind, which is 31 years (BfS:2024).
-The sample is further refined to include only women who gave birth to their first child between 20 and 40.111111By construction, the oldest women in the sample are 37 years old when they have their first child - women born 1985 that have their first child in 2022.
+The sample is further refined to include only women who gave birth to their first child between 20 and 40.121212By construction, the oldest women in the sample are 37 years old when they have their first child - women born 1985 that have their first child in 2022.
 Hence, only the lower age is binding.
-This age range allows us to focus on women’s reproductive years.
+This age range allows us to focus on womens’ reproductive years.
 We further constrain the samples to the years between 2012 and 2022, a choice influenced by the introduction of the DRG on the 1st of January 2012 (DRG:2024).
 [^13] A Diagnostic-Related Group (DRG) is a system used by health insurance companies to categorise hospitalisation costs and determine payment amounts for hospital stays.
-Instead of billing for each individual service received, a fixed payment amount is predetermined based on the DRG assigned to your stay..
+Instead of billing for each individual service received, a fixed payment amount is predetermined based on the DRG assigned to your stay.
 This significant event in the healthcare system could affect data reliability prior to 2012.
@@
 This would reduce the sample size even more.
+Specifically, individuals with poorer mental health may have less energy or motivation to change their insurance provider, leading to a sample that overrepresents those with mental health challenges.
 Table 3.3.2 in Section 3.3.2 shows that the balanced sample only differs from the unbalanced sample in terms of nationality, language, and number of children.
@@
 Hence, the power of the analysis is reduced.
-An unbalanced sample is used to check for robustness.
+An unbalanced sample is used to check for robustness due to the reasons mentioned in the previous paragraph.
 Furthermore, a balanced sample only going until 2019 ($N=18,714$) is used to avoid the potential bias of the COVID-19 pandemic on mothers’ mental health.
@@
 [^16] A robustness check with the non-smoothed outcome can be found in Appendix LABEL:appendix_non_smoothed.
-Nevertheless, antidepressant prescriptions are potentially still a noisy proxy for mental health, and they most likely underestimate mental health problems, as they are only prescribed in severe cases.
+Nevertheless, antidepressant prescriptions are potentially still a noisy measure of mental health, and they most likely underestimate mental health problems, as not everyone with mental health issues receives medication.
+Antidepressants are typically prescribed for moderate to severe cases, but many individuals with milder symptoms rely on therapy, other treatments, or do not seek treatment at all (Obsan:2022).
 In addition to the primary outcome, we consider other mental health proxies, such as the number of visits to a GP or psychiatrist.
@@
 Therefore, there is a slight measurement error in this outcome variable.
-Furthermore, the number of visits to the GP or HMO can only be identified consistently from 2012 to 2022 for the CSS data, which leads to a smaller sample size for this outcome.
+Moreover, the number of visits to the GP or HMO can only be identified consistently from 2012 to 2022 for the CSS data, which leads to a smaller sample size for this outcome.
 The outcome variables are measured relative to the event-time of birth $t$, with $t+1,t+2,\dots$ representing the months post-childbirth, and $t-1,t-2,\dots$ the months pre-childbirth.
@@
 The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval.
+The slow decrease in antidepressants at the start of the pregnancy and also the slow increase after giving birth are partly coming from the smoothing of the antidepressant prescriptions.
+Compared to Denmark or Austria, the antidepressant prescription rate is lower before childbirth.
+A reason, therefore, could be that even after monthly health insurance payments, Swiss individuals have to pay for medical visits and drugs up to a yearly deductible (often CHF 2,500).
 ($N=25,120$)
@@
 Since the prescription data only begins in 2012, identifying actual first-time prescriptions is challenging.
-As we move further in time, the likelihood increases that the observed prescriptions are genuinely first-time prescriptions, explaining the negative trend in antidepressant use among women without children (blue line) in Figure 2.
-To minimise the problem, we select women with at least one year without antidepressant prescriptions before their first recorded prescription.
-The number of women receiving their first antidepressant prescription stabilises at a consistent level in the years after childbirth.
-This suggests that the observed increase in antidepressant use is not only explainable by postpartum depression.
+As we move further in time, the likelihood increases that the observed prescriptions are genuinely first-time prescriptions, explaining the negative trend in antidepressant use among women without children (blue line) in Figure 2.181818For the women without children placebo dates for the first birth are randomly drawn.
+We follow Ahammer:2023 and approximate the factual distribution of age at first birth by a birth-cohort specific log-normal distribution $LN(\mu,\sigma^{2})$.
+The month is randomly drawn from a uniform distribution.
+To minimise this issue, we select women who had at least one year without antidepressant prescriptions before their first recorded prescription.
+We find that the number of women receiving their first antidepressant prescription stabilizes at a consistent level in the years following childbirth.
+This suggests that the observed increase in antidepressant use is not solely driven by postpartum depression but persists even two to six years after childbirth, indicating a longer-term effect.
 Figure 2: Descriptives: First-time antidepressants prescriptions
@@
 The blue line shows the linear trend for women who never have a child but get assigned random birthdates.
+The factual distribution of age at first birth for women without children has been approximated by a birth-cohort specific log-normal distribution $LN(\mu,\sigma^{2})$.
+The month is randomly drawn from a uniform distribution.
 The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval.
@@
 Controlling for age is important, as the risk of depression increases with age and because women who are mentally unhealthy might postpone having a child.
-Hence, women having a child at different ages may have different mental health trends.
+Consequently, women giving birth at different ages may exhibit different underlying trends in mental health, which could bias estimates if not accounted for.
 The average age of having a first child is around 29 years, and the age distribution is shown in Figure 5.202020The average age of Swiss women when they give birth for the first time is 30.
@@
 Note: The figure shows the age distribution of women having a child between 2012 and 2022. ($N=25,120$)
-Accident coverage is a proxy for employment status, as it is typically needed by those unemployed, out of the labour force, or self-employed.
-At the same time, employers generally provide it for their employees.
-Figure 6 shows a notable increase in accident coverage following childbirth, suggesting a rise in non-employment among women.
-Around 80% of women do not have accident coverage before childbirth, implying an 80% employment rate.
-[^19] The decrease in accident coverage pre-birth is partly due to young women and men not yet having accident insurance because they are, for example, studying or doing an apprenticeship.
+Employment status can be identified by looking at the accident coverage of the health insurance.
+Employers generally provide it for their employees, hence, employed individuals do not have accident coverage in the health insurance.
+On the other hand, it is needed by those unemployed, out of the labour force or self-employed.
+Figure 6 shows a notable decrease in employement following childbirth.
+Around 80% of women are employed before childbirth.
+[^21] The increase in employment pre-birth is partly due to young women and men not yet started to work as they still are, for example, studying or doing an apprenticeship.
 This contrasts with the 90% pre-childbirth employment rate for women in Switzerland as reported by official statistics, indicating a discrepancy due to self-employment (BfS:2022).
 8.1% of women without and 11.8% of women with children in Switzerland are self-employed.
-As depicted by Figure 6, accident coverage rises to about 30% (70% employed) post-childbirth, and further increases to 40%, suggesting 60% employment.
+As depicted by Figure 6, employment rates drop to about 70% post-childbirth, and further reduce to 60%.
 Official figures show a decrease in employment to 80% after the first child and 69% after the second (BfS:2022).
@@
 Since accident insurance is mandatory and reminders are sent if it is not in place, this issue is less pronounced.
-Figure 6: Descriptives: Accident coverage
-Note: The figure shows the proportion of women having accident coverage around the time of birth ($t=0$: time of first birth).
+Figure 6: Descriptives: Employment
+Note: The figure shows the proportion of women employed around the time of birth ($t=0$: time of first birth).
 The red shaded area illustrates the pregnancy, and the shaded area represents the 95% confidence interval.
 ($N=25,120$)
-The left part of Table 3.3.2 shows the sociodemographic characteristics of women in the month they have a child, separated by the different samples used, whereas the right part of the table shows it separated by women having an antidepressant prescription or not.
+The left part of Table 3.3.2 presents the sociodemographic characteristics of women in the month they give birth, across the different samples used in the analysis.
+This helps to evaluate whether the sample selection might introduce bias and whether robustness checks are warranted.
+The right part of the table compares women with and without an antidepressant prescription prior to childbirth, highlighting potentially meaningful differences between these groups that motivate further investigations of heterogeneous treatment effects.
 In the balanced sample, 19% of women receive subsidies for health care, 63% of women have more than one child, and 30% have a cesarean section.
@@
 By construction, the oldest woman giving birth can only be 34 years old instead of 37, which pushes the average age downwards.
-The unbalanced has more women with only one child by construction.
+The unbalanced sample has more women with only one child by construction.
 We do not see all the years for the women in the unbalanced sample, hence, we miss some second births.
-Comparing women who have an antidepressant prescription with those who do not, we see that slightly more women speaking French and fewer women speaking German have an antidepressant prescription.
-This pattern can also be seen in the regions of residence, as more women living in the French-speaking part of Switzerland have an antidepressant prescription compared to, for example, central Switzerland.
-As expected, women with an antidepressant prescription are more likely to have an accident insurance, meaning that they are more likely to not be employed.
-Furthermore, they are more likely to have a subsidy for health care, have a cesarean section and only have one child.
+As there are some differences between the two samples, the unbalanced sample is used as a robustness check (see Section LABEL:robustness:unbalanced_sample).
 Table 1: Descriptives: Comparison of different samples for staggered difference-in-difference analysis